### v7

@@ 1 Introduction and overview
Some other economists, including John Rae (1834) and John Stuart Mill (1848) (see Appendix C), argued that capital growth might also be explained by a rise in productivity of capital and labor already extant. Ways might found for existing factors to produce more, that is, and so to allow more consumption, or more capital growth, or any mix of the two, without inputs of net saving. Robert Solow (1957) allowed that possibility for “disembodied” growth, where plant and products already existing are repurposed or redeployed in more productive ways.
We test between those two explanations of capital growth, by net saving or by increase in productivity of capital and labor already in existence, by comparing net saving to concurrent change in market-value capital in 86 countries. As changes in net saving are expected to be associated with opposite changes in consumption, we also compare change in consumption to concurrent change in capital growth (capital acceleration). All data are drawn from national accounts of those countries as collated on the free website World Inequality Database.
Those test results from data for net saving/capital ratio and for consumption/capital ratio are easiest to reconcile in an interpretation that change in neither affects change in capital growth rate. [^2] Another interpretation allows that net saving may account for 6.4% of capital growth. See Section 4 below. This finding would support the views of Rae and Mill, and of Solow as to disembodied growth. It would suggest that capital growth, even in acceleration, arrives without help from net saving or consumption restraint. Net saving, if so, raises the physical quantity of capital, but not the aggregate value, and so reduces the value per unit.
@@
It predicts only at the largest scales, and only for the private sector. Individuals and groups and even small economies can grow through investment from outside. That possibility is foreclosed only at the scale of all capital and all economies together. The public sector, meanwhile, responds to political rather than market choices, and grows or shrinks accordingly.
@@ 2 Net saving, capital growth and capital acceleration
This study compares net saving $S_{net}$ to concurrent growth in market-value capital $K$ from data in national accounts. Capital growth ${\Delta K}_{i}$ in each year $i$ for each reporting country is found as $K_{i}-K_{i-1},$ and compared to $S_{net,i}$ reported for that year and country. As we will be testing for differences between $\Delta K$ and $S_{net,i}$, we begin by writing
$\Delta K_{i}=z_{i}S_{net,i}+Q_{i}\ .$ (1)
Here "net saving efficiency" $z$ is the fraction of $S_{net}$ actually realized in growth of capital value $K$. $Q$ means the sum of market noise, which may prove positive or negative or zero, plus any part of capital growth explained by concurrent productivity gain as described by Rae and Mill, and by Solow as to disembodied growth. We call this sum of noise and productivity gain “free growth” in that it costs no net saving.
The object of testing is to find effects of $S_{net}$ on concurrent $\Delta K$, and so to help evaluate historic and current teachings as to those effects. We submitted above that most teaching, with exceptions noted as to Rae, Mill, and Solow, and within the warranted rate, predicts net saving to differ from concurrent growth in market-value capital only by market noise which tends to balance out over scale and time. Free growth $Q$ in Eq. (1), in that case, will give the market noise converging to zero. That consensus prediction, which we will challenge, will here be called “thrift theory”; $S_{net}$, in thrift theory, is expected to converge to $\Delta K$ if held within the warranted rate. That is,
$E(z)=1\ ,\quad E(\Delta K)=S_{net}\quad\text{and}\quad E(Q)=0,\quad\text{if}\quad\frac{\Sigma S_{net}}{\Sigma K}\leq u\ ,\quad\text{in thrift theory,}$, (2)
where:
  1. 1.
    $\Sigma S_{net}$ is collective net saving over the economy
  2. 2.
    $\Sigma K$ is collective capital over the economy before current $\Sigma S_{net},$ and
  3. 3.
    $u$ is the warranted rate.
Free growth theory makes the opposite prediction that net saving adds no capital value, and that capital growth, at the largest scales in the private sector, is explained wholly by increased productivity of skills and materials already present. Thus, referring to Eq. (1), it predicts:
$E(z)=0\quad\text{and}\quad E(\Delta K)=Q\quad\text{in free growth theory.}$. (3)
$E(z)$, $E(\Delta K)$ and $E(Q)$ here give the expected values of $z$, $\Delta K$ and $Q$ respectively. Expected value means predicted average of outcomes over all observations. We will test those predictions against data for saving and capital growth taken from national accounts. As secular economic growth has tended to make later stocks and flows larger than earlier ones, we first divide by $K$ (normalize) to avoid overweighting of more recent years in finding that average. Division of Eq. (1) by $K_{i-1}$ gives
$\frac{(\Delta K)_{i}}{K_{i-1}}=\frac{z_{i}S_{net,i}}{K_{i-1}}+\frac{Q_{i}}{K_{i-1}}\ .$ (4)
The first term in Eq. (4) gives capital growth rate $g(K)$. The fraction in the second term is a variant of the Keynesian net saving rate $s_{net}$ where capital rather than output becomes the denominator. This flow will show as $s^{*}$, with the subscript “net” left implicit, and with the understanding that the denominator shows market-value capital rather than output. The third term in Eq. (4) will be called free growth rate and shown as $q_{s}$. That is,
$g{(K)}_{i}=\frac{(\Delta K)_{i}}{K_{i-1}}\ ,\quad s^{*}_{i}=\frac{S_{net,i}}{K_{i-1}}\ ,\quad\text{and}\quad q_{i}=\frac{Q_{i}}{K_{i-1}}\ ,$
so that Eq. (4) can be shown more compactly as
$g(K)_{i}=z_{i}s^{*}_{i}+q_{i}\ .$ (5)
By the definition $q=\frac{Q}{K}$, an expected value $E(Q)=0$ implies $E(q)=0$. Application of Eq. (2) to Eq. (5) now gives
$E(g(K))=s^{*}\quad\text{and}\quad E(q)\ ,\quad\text{under thrift assumptions,}$, (6)
where “thrift assumptions” are that thrift theory is correct and that the warranted rate is not exceeded. Eqs. (2) and (4), meanwhile, give
$E(g(K))=q\quad\text{in free growth theory.}$. (7)
Note here that free growth theory makes no prediction as to the quantity $s^{*}_{i}$. Rather it predicts $E(z)=0$ in Eq. (5), so that $s^{*}_{i}$ takes no part in calculation of $\Delta g(K)_{i}$.
The prediction in Eq. (6) can be tested directly from data for market-value capital and net saving in national accounts. We define the "auxiliary thrift index" by
θ=sg(K),that is, and test the prediction
@@
[…]
$E(\theta^{*})=1\quad\text{under thrift assumptions.}$.
Results show in Fig. 1 and Tables 1 and 2. They show $\theta^{*}$ as averaging something near $\frac{1}{2}$, and with wide variance from country to country and from year to year. Regression of $s^{*}$ on $g(K)$ is found at 0.2103. Free growth theory, again, makes no prediction as to relationships between net saving and capital growth.
The concepts of correlation, covariance and regression are founded on the principle that causal relationships among variables tend to be revealed in simultaneous changes. Thus the two theories are best tested in their implications for changes in capital growth, which we call capital acceleration and show as $\Delta g(K)$. First note that Eq. (5) allows
$(\Delta g(K))_{i}=(\Delta(zs^{*}))_{i}+(\Delta g)_{i}\ ,\quad\text{where}$ (8)
$(\Delta g{(K)})_{i}=(g{(K)})_{i}-(g(K))_{i-1}\ ,\quad(\Delta(zs^{*}))_{i}=z_{i}s^{*}_{i}-z_{i-1}s^{*}_{i-1}\ ,\quad\text{and}\quad(\Delta q)_{i}=q_{i}-q_{i-1}\ .$
For any variables $a$ and $b$, we may reason $E(a-b)=E(a)-E(b)$. By this and by Eqs. (7) and (8), then,
$E(g(K))=\Delta s^{*}\quad\text{and}\quad E(\Delta q)=0\ ,\quad\text{under thrift assumptions.}$. (9)
Division of Eq. (8) by capital acceleration, and rearrangement, gives
$\frac{(\Delta(zs^{*}))_{i}}{(\Delta g{(K)})_{i}}+\frac{(\Delta q)_{i}}{(\Delta g{(K)})_{i}}=1\ .$ (10)
Fig. 1: $\theta^{*}$ (86 countries)
Table 1: Regression of $s^{*}$ on $g(K)$ (Screen = 0.01). $H_{0}$ per thrift theory: $g(K)\cong s^{*}$ $\overline{\theta^{*}}\cong 1$.
$s^{*}$
Regression of $s^{*}$ on $g(K)$ 0.1857∗∗∗
@@
Define $\theta_{s,i}=\frac{(\Delta zs^{*})_{i}}{(\Delta g{(K)})_{i}}$ and $\varphi_{s,i}=\frac{(\Delta q)_{i}}{(\Delta g{(K)})_{i}}$ to restate Eq. (10) as
$\theta_{s,i}+\varphi_{s,i}=1\ .$ (11)
Next define $\theta_{s}=E(\theta_{s,i})$ and $\varphi_{s}=E(\varphi_{s,i})$. By Eq. (11), then,
[…]
$\theta_{s}=1\quad\text{and}\quad\varphi_{s}=0\ ,\quad\text{under thrift assumptions.}$. (13)
Eq. (11) enables measurement of effects of thrift and free growth in a given period $i$, that is, while Eq. (12) does the same for all periods together.
$\theta$ and $\varphi$ will be called the “thrift index” and “free growth index” respectively. $\theta_{s}$ and $\varphi_{s}$ give their values as found from data for change in net saving. Expected values, again, are predicted averages of outcomes. Thus Eq. (13) and thrift theory can be tested by finding average values of $\theta_{s,i}$ and comparing findings to the expected value $\theta=1$. First we find
$\overline{\theta_{s,i}}=\frac{1}{m}\sum^{m}\theta_{s,i}\quad\text{and}\quad\overline{\varphi_{s,i}}=\frac{1}{m}\sum^{m}\varphi_{s,i}\ ,$
@@
Table 2: Average $\theta^{*}$ in 86 countries (screen = 0.01). Number of years clearing screen shown in ()
@@
Country Period $\overline{\theta^{*}}$ Armenia 1996 - 2020 (23) -0.63 Aruba 1996 - 2001 (6) 1.25 Australia 1980 - 2018 (33) 0.24 Austria 1996 - 2021 (23) 0.83 Azerbaijan 1996 - 2020 (23) 1.57 Bahrain 2009 - 2013 (5) -2.43 Belgium 1996 - 2021 (21) 0.61 Bolivia 1998 - 2015 (17) 0.40 Botswana 1996 - 1999 (4) 3.54 Brazil 1996 - 2019 (21) 0.25 British Virgin Islands 1996 - 1999 (4) 2.45 Bulgaria 1996 - 2016 (16) 0.14 Burkina Faso 2000 - 2019 (20) 0.23 Cameroon 1996 - 2019 (24) 1.10 Canada 1980 - 2022 (38) 0.37 Cape Verde 2008 - 2017 (9) 0.53 Chile 1997 - 2021 (19) 0.84 China 1992 - 2020 (29) 0.81 Colombia 1996 - 2022 (27) 0.60 Costa Rica 2013 - 2020 (6) 0.21 Croatia 1996 - 2021 (24) 0.43 Curaçao 2001 - 2016 (14) 1.26 Cyprus 1996 - 2021 (25) 0.07 Czechia 1994 - 2020 (20) 0.17 Côte d’Ivoire 1996 - 2000 (5) 0.29 Denmark 1996 - 2022 (26) 0.22 Dominican Republic 1996 - 2016 (10) 1.52 Ecuador 2008 - 2020 (13) 0.53 Egypt 1997 - 2015 (19) 0.92 El Salvador 2015 - 2019 (5) -0.62 Estonia 1996 - 2021 (24) 0.45 Finland 1996 - 2021 (23) 0.17 France 1980 - 2021 (35) 0.26 Germany 1980 - 2022 (38) 0.70 Greece 1996 - 2021 (26) 0.17 Guatemala 2002 - 2021 (20) -0.72 Guinea 2004 - 2010 (5) 0.72 Honduras 2001 - 2015 (14) -0.57 Hong Kong 1997 - 2011 (15) 0.90 Hungary 1996 - 2021 (21) 0.14 Iceland 2001 - 2014 (14) 0.22 India 1999 - 2019 (21) 0.57 Indonesia 2017 - 2019 (3) 0.94 Iran 1997 - 2017 (20) 4.96 Ireland 1996 - 2021 (22) 0.26 Israel 1996 - 2019 (24) 0.46 Country Period $\overline{\theta^{*}}$ Italy 1980 - 2022 (36) 0.13 Japan 1980 - 2021 (29) 0.19 Kazakhstan 1997 - 2022 (25) 0.60 Kuwait 2003 - 2017 (14) 1.11 Kyrgyzstan 1996 - 2021 (24) 0.46 Latvia 1996 - 2021 (24) -0.23 Lithuania 1996 - 2021 (24) 0.13 Luxembourg 1996 - 2021 (22) 0.03 Malaysia 2007 - 2015 (9) 2.52 Malta 1996 - 2021 (23) 0.16 Mauritius 2014 - 2018 (5) 0.04 Mexico 1996 - 2021 (22) 0.15 Moldova 1996 - 2019 (22) -0.91 Mongolia 2006 - 2020 (15) 0.10 Morocco 1999 - 2021 (23) 1.13 Netherlands 1996 - 2021 (23) 0.51 New Zealand 1996 - 2019 (21) 0.78 Nicaragua 2007 - 2018 (12) 0.08 Niger 1996 - 2019 (24) 0.90 Norway 1981 - 2021 (38) 0.95 Peru 2008 - 2021 (14) 0.76 Philippines 1996 - 2022 (27) 1.76 Poland 1996 - 2021 (25) 0.71 Portugal 1996 - 2022 (24) 0.00 Qatar 2002 - 2017 (14) 1.89 Romania 1996 - 2020 (22) 0.03 Russia 1996 - 2017 (13) -0.19 Saudi Arabia 2003 - 2009 (7) 2.65 Senegal 2015 - 2021 (7) 0.24 Serbia 1998 - 2021 (23) 0.61 Slovakia 1996 - 2022 (24) 0.15 Slovenia 1996 - 2021 (20) 0.28 South Africa 1996 - 2022 (24) 0.28 South Korea 1996 - 2020 (24) 0.63 Spain 1995 - 2021 (25) 0.32 Sweden 1980 - 2021 (37) 0.41 Switzerland 1993 - 2021 (25) 0.69 Tunisia 1996 - 2011 (16) 0.29 Turkey 2010 - 2017 (8) 1.32 USA 1980 - 2021 (38) 0.31 Ukraine 1996 - 2019 (21) 0.17 United Kingdom 1981 - 2021 (36) 0.16 Uruguay 2016 - 2017 (2) 1.16 Uzbekistan 2011 - 2021 (10) 0.21 Vanuatu 2002 - 2007 (6) 0.29 Venezuela 1998 - 2019 (20) 0.55
@@
Note: Thrift theory predicts $\overline{\theta^{*}}\cong 1$. Free growth theory makes no prediction for these data.
where $m$ is the number of observed values of $\theta_{s,i}$ and $\varphi_{s,i}$, and test the predictions
$\overline{\theta_{s,i}}\cong 1\quad\text{and}\quad\overline{\varphi_{s,i}}\cong 0\ ,\quad\text{under thrift assumptions.}$.
Calculations of $\overline{\theta_{s,i}}$ and $\overline{\varphi_{s,i}}$ are not expected to show 1 and 0 exactly, under thrift assumptions, so that $\overline{\theta_{s,i}}$ and $\overline{\varphi_{s,i}}$ are not exactly equal to $\theta_{s}$ and $\varphi_{s}$, because the number of samples $m$ is finite.
Free growth theory, maenwhile, referring to Eq. (13), does not predict the exact opposite findings $\theta_{s}=0$ and $\varphi_{s}=1$. It rather expects a value of $\varphi_{s}$ somewhat less than unity for reasons shown in Section 4.
Fig. 2 shows average values of $\theta_{s,i}$ and $\varphi_{s,i}$ for 86 countries, both unweighted and weighted to GDP, over the period 1980-2022. Table 3 shows associated regressions of $\Delta s^{*}$ and $\Delta g_{s}$ respectively on $\Delta g(K)$ at 0.1000 and 0.9000. Average values of $\varphi_{s}$ for the 86 individual countries show in Table 4. To control distortions brought by small absolute denominators, years were screened out where $|\Delta g(K)|$ was found at less than 0.01 (see Section 13). Results show $\overline{\varphi_{s,i}}=0.936$ and $\overline{\theta_{s,i}}=0.064$ for data weighted to GDP. These findings largely support free growth theory, but could suggest a small role (6.4%) for thrift as well. This possibility will be discussed in Section 4. We will first test thrift theory from a different approach.
Fig. 2: Thrift and free growth indexes derived from net saving (86 countries).
Table 3: Regression of $\Delta s^{*}$ and $\Delta q_{s}$ on $\Delta g(K)$ (Screen = 0.01). $H_{0}\ \text{per thrift theory:}\ \Delta g(K)\cong\Delta s^{*}\ \&\ \Delta q_{s}\cong 0$
$\Delta s^{*}$ $\Delta q_{s}$
Regression of value shown on $\Delta g(K)$ 0.1000∗∗∗ 0.9000∗∗∗
(0.0145) (0.0145)
Observations 1,574 1,574
R2 0.32867 0.96233
Within R2 0.22713 0.95965
Year fixed effects $\checkmark$ $\checkmark$
Country fixed effects $\checkmark$ $\checkmark$
3 Consumption and capital growth
The concept of cash flow is little used in economics, aside from finance and in present value calculations. It is defined implicitly in the equation
$P=\Delta K+F\ ,$ (14)
where $P$ and $F$ mean net profit and cash flow respectively [^3] We have found no standard notation for cash flow.. Cash flow means the net outflow of value, in cash or kind, from an asset to its owner for consumption $C$ or for reinvestment in other assets. This flow of spending is augmented by pay $\Pi$ [^4] Compensation of employees plus a share of mixed income. We avoid the standard notation $W$, which is also used for work, as we prefer to distinguish pay from work.. At the scale of the closed economy, meaning of all assets together, reinvestment cancels to zero. At this scale, then, the sum of spending $F+\Pi$ is spent on consumption alone. That is,
$\Pi+F=C$ (15)
at the collective scale. Eq. (14), meanwhile, may be arranged as $F=P-\Delta K$. Substitution of this in Eq. (15), with rearrangement, gives
$\Delta K=\Pi+P-C\ .$ (16)
Thrift theory, which expects $\Delta K=S_{net}$, writes this as
$S_{net}=\Pi+P-C\ ,\quad\text{under thrift assumptions.}$. (17)
Pay, net profit [^5] Jordà et al. (2019), The rate of return on everything. and consumption can all be measured, to an extent, as can market-value capital, so that this prediction of thrift theory could be measured directly. Comparison between Eqs. (16) and (17), however, shows that such a test, given perfect data in national accounts, would only repeat that of Eq. (6). As before, the two theories are tested most revealingly in capital acceleration. Given the definitions [^6] The notation $c^{*}$, as with $s^{*}$, shows an asterisk to distinguish its meaning $c^{*}=\frac{C}{K}$ from the Keynesean $c=\frac{C}{Y}$
$\pi_{i}=\frac{\Pi_{i}}{K_{i-1}}\ ,\quad p_{i}=\frac{P_{i}}{K_{i-1}}\quad\text{and}\quad c^{*}_{i}=\frac{C_{i}}{K_{i-1}}\ ,$
Eq. (16) can be divided by capital to reach
$g(K)_{i}=\pi_{i}+p_{i}-c^{*}_{i}\ ,\quad\text{implying}$ (18)
$(\Delta g(K))_{i}=\Delta\pi_{i}+\Delta p_{i}-\Delta c^{*}_{i}\ ,\quad\text{where}$ (19)
$\Delta\pi_{i}=\pi_{i}-\pi_{i-1}\ ,\quad\Delta p_{i}=p_{i}-p_{i-1}\quad\text{and}\quad\Delta c^{*}_{i}=c^{*}_{i}-c^{*}_{i-1}\ ,$
as before. Thrift theory expects $\Delta g(K)=\Delta s^{*}=\Delta\pi+\Delta p-\Delta c$ by Eqs. (9) and (17). It further expects changes in productivity, as realized in the pay/capital and profit/capital ratios $\pi$ and $p$, to converge to zero as market noise, so that the only source of net saving is in consumption restraint $-\Delta c^{*}$. Thus
$E(\Delta g(k))=-\Delta c^{*}\quad\text{and}\quad E(\Delta\pi)=E(\Delta p)=0\ ,\quad\text{under thrift assumptions.}$. (20)
Free growth theory, conversely, predicts that net saving contributes nothing to capital growth, and hence that net saving through consumption restraint, at the collective scale and in the private sector, is sacrifice with nothing in return. Free growth theory expects rather that costless increases in the ratios of pay and net profit to capital, enabled by greater productivity of human and physical capital already present, account for all capital growth and acceleration. Referring to Eq. (19), however, it predicts
[…]
for reasons shown in Section 4. As in Eq. (10), we prepare to test thrift theory by first dividing Eq. (20) by capital acceleration to reach
$E(\frac{-\Delta c^{*}}{\Delta g(K)})=1\ ,\quad\text{under thrift assumptions.}$. (22)
Define $\theta_{c,i}=\frac{-\Delta c^{*}_{i}}{\Delta g(K)_{i}}$ and $\varphi_{c,i}=1-\theta_{c,i}$ to re-express Eq. (22) as
$\theta_{c,i}+\varphi_{c,i}=1\ .$ (23)
By Eqs. (20) through (23), further,
$E\left(\theta_{c}\right)=1\quad\text{and}\quad E\left(\varphi_{c}\right)=0\ ,\quad\text{under thrift assumptions.}$. (24)
Define $\theta_{c}=E\left(\theta_{c}\right)$ and $\varphi_{c}=E\left(\varphi_{c}\right)$ to re-express Eqs. (23) and (24) respectively as
$\theta_{c}+\varphi_{c}=1\ ,\quad\text{and}$ (25)
$\theta_{c}=1\quad\text{and}\quad\varphi_{c}=0\ ,\quad\text{under thrift assumptions.}$. (26)
We infer $\theta_{c}\cong\overline{\theta_{c,i}}$ and $\varphi_{c}\cong\overline{\varphi_{c,i}}$ as before, and test thrift theory by comparing average yearly values of $\theta_{c,i}$ and $\varphi_{c,i}$ to its predictions $\overline{\theta_{c,i}}\cong 1$ and $\overline{\varphi_{c,i}}\cong 0$.
[…]
Fig. 3: Thrift and free growth indexes derived from consumption (86 countries).
@@ 3 Consumption and capital growth
Table 4: Average $\varphi_{s,i}$ and $\varphi_{c,i}$ in 86 countries (screen = 0.01). Number of years clearing screen shown in ()

### v8

@@ 1 Introduction and overview
Some other economists, including John Rae (1834) and John Stuart Mill (1848) (see Appendix C), argued that capital growth might also be explained by a rise in productivity of capital and labor already extant. Ways might be found for existing factors to produce more, that is, and so to allow more consumption, or more capital growth, or any mix of the two, without inputs of net saving. Robert Solow (1957) allowed that possibility for “disembodied” growth, where plant and products already existing are repurposed or redeployed in more productive ways.
We test between those two explanations of capital growth, by net saving or by increase in productivity of capital and labor already in existence, by comparing net saving to concurrent change in market-value capital in 92 countries. All data are drawn from national accounts of those countries as collated on the free website World Inequality Database.
Those test results from data for net saving/capital ratio are easiest to reconcile in an interpretation that net saving has no effect on capital growth. Net saving, if so, raises the physical quantity of capital, but not the aggregate value, and so reduces the value per unit.
@@
It predicts only at the largest scales, and only for the private sector. Individuals and groups and even small economies can grow or decline through net saving or dissaving. That possibility is foreclosed only at the scale of all capital and all economies together. The public sector, meanwhile, responds to political rather than market choices, and grows or shrinks accordingly.
@@ 2 Net saving, capital growth and capital acceleration
Thrift theory will mean the combined ideas that all net saving $S_{net}$ is realized in equal growth of capital $K$, if $S_{net}$ holds within the warranted rate, and that there is no other source of capital growth except market noise regressing to zero. "Thrift assumptions" will mean those two plus the assumptions that current $S_{net}$ holds within the warranted rate. The acronym uta will mean "under thrift assumptions." Then
$\Delta K=S_{net}\ ,\quad\text{uta}.$ (1)
Define "thrift" $s^{*}$ by $s^{*}=\frac{S_{net}}{K}$, and divide Eq. (1) by capital to predict capital growth rate $g(K)$ as
$g(K)=s^{*}\ ,\quad\text{uta,}\quad\text{where }s^{*}=\frac{S_{net}}{K}.\quad\text{ Then}$ (2)
@@
$\Delta g(K)$ and $\Delta s^{*}$ respectively will be called "capital acceleration" and "thrift change," either of which may be positive or negative. Division of Eq. (3) by capital acceleration, with rearrangement, gives
$\frac{\Delta s^{*}}{\Delta g(K)}=\frac{\Delta g(K)}{\Delta g(K)}=1\ ,\quad\text{uta}.$ (4)
The argument has continued from growth in Eq. (2) to acceleration in Eqs. (3) and (4) on the reasoning that causal relationships among variables, in this case between thrift and capital growth rate, tend to be revealed most reliably in their concurrent changes. [^3] The ratio $\frac{\Delta A}{\Delta B}$, that is, compares changes in A to concurrent changes in B, and so suggests a causal relationship between A and B insofar as that ratio holds constant. This is the principle on which the concepts of correlation, covariance and regression are founded. Eq. (4) was derived from Eq. (3) so that success of predictions from growth assumptions can be measured against a standard of unity. For notational convenience, then, we define the "thrift index" $\theta$ by $\theta=\frac{\Delta s^{*}}{\Delta g(K)}$, and restate Eq. (4) as
$\theta=1\ ,\quad\text{uta,}\quad\text{where}\ \theta=\frac{\Delta s^{*}}{\Delta g(K)}\ .$ (5)
Test results for the predictions $\frac{s^{*}}{g(K)}=\frac{\Delta s^{*}}{\Delta g(K)}=1$, under thrift assumptions, appear in Fig. 1 and in Tables 1 and 2. Results were found as GDP-weighted averages over all countries and years, from data for net saving and market-value capital taken from national accounts. They show $\frac{s^{*}}{g(K)}=0.481$, with regression of $s^{*}$ on $g(K)$ found at 0.0771, and $\frac{\Delta s^{*}}{\Delta g(K)}=0.064$, with regression of $\Delta s^{*}$ on $\Delta g(K)$ at 0.0559.
Fig. 1: Average $\frac{s^{*}}{g(K)}$ with LOESS smoothing over all countries, GDP-weighted, 1980-2022
Table 1: Regression of $s^{*}$ on $g(K)$ and $\Delta s^{*}$ on $\Delta g(K)$, GDP-weighted. Screen = 0.01. $H_{0}$ per thrift theory: $\frac{s^{*}}{g(K)}=\frac{\Delta s^{*}}{\Delta g(K)}=1$.
Regression of $s^{*}$ on $g(K)$ 0.0771∗∗∗
@@
Regression of $\Delta s^{*}$ on $\Delta g(K)$ 0.0559∗∗∗
Observations 1,574
R2 0.37071
Within R2 0.27233
Year fixed effects $\checkmark$
Country fixed effects $\checkmark$
@@
Table 2: Average $\frac{s^{*}}{g(K)}$ in 92 countries (screen = 0.01). Number of years clearing screen shown in ()
@@
Armenia 1996 - 2020 (23) -0.63 Aruba 1996 - 2001 (6) 1.25 Australia 1980 - 2018 (33) 0.24 Austria 1996 - 2021 (23) 0.83 Azerbaijan 1996 - 2020 (23) 1.57 Bahrain 2009 - 2013 (5) -2.43 Belgium 1996 - 2021 (21) 0.61 Bolivia 1998 - 2015 (17) 0.40 Botswana 1996 - 1999 (4) 3.54 Brazil 1996 - 2019 (21) 0.25 British Virgin Islands 1996 - 1999 (4) 2.45 Bulgaria 1996 - 2016 (16) 0.14 Burkina Faso 2000 - 2019 (20) 0.23 Cameroon 1996 - 2019 (24) 1.10 Canada 1980 - 2022 (38) 0.37 Cape Verde 2008 - 2017 (9) 0.53 Chile 1997 - 2021 (19) 0.84 China 1992 - 2020 (29) 0.81 Colombia 1996 - 2022 (27) 0.60 Costa Rica 2013 - 2020 (6) 0.21 Croatia 1996 - 2021 (24) 0.43 Curaçao 2001 - 2016 (14) 1.26 Cyprus 1996 - 2021 (25) 0.07 Czechia 1994 - 2020 (20) 0.17 Côte d’Ivoire 1996 - 2000 (5) 0.29 Denmark 1996 - 2022 (26) 0.22 Dominican Republic 1996 - 2016 (10) 1.52 Ecuador 2008 - 2020 (13) 0.53 Egypt 1997 - 2015 (19) 0.92 El Salvador 2015 - 2019 (5) -0.62 Estonia 1996 - 2021 (24) 0.45 Finland 1996 - 2021 (23) 0.17 France 1980 - 2021 (35) 0.26 Germany 1980 - 2022 (38) 0.70 Greece 1996 - 2021 (26) 0.17 Guatemala 2002 - 2021 (20) -0.72 Guinea 2004 - 2010 (5) 0.72 Honduras 2001 - 2015 (14) -0.57 Hong Kong 1997 - 2011 (15) 0.90 Hungary 1996 - 2021 (21) 0.14 Iceland 2001 - 2014 (14) 0.22 India 1999 - 2019 (21) 0.57 Indonesia 2017 - 2019 (3) 0.94 Iran 1997 - 2017 (20) 4.96 Ireland 1996 - 2021 (22) 0.26 Israel 1996 - 2019 (24) 0.46 Italy 1980 - 2022 (36) 0.13 Japan 1980 - 2021 (29) 0.19 Kazakhstan 1997 - 2022 (25) 0.60 Kuwait 2003 - 2017 (14) 1.11 Kyrgyzstan 1996 - 2021 (24) 0.46 Latvia 1996 - 2021 (24) -0.23 Lithuania 1996 - 2021 (24) 0.13 Luxembourg 1996 - 2021 (22) 0.03 Malaysia 2007 - 2015 (9) 2.52 Malta 1996 - 2021 (23) 0.16 Mauritius 2014 - 2018 (5) 0.04 Mexico 1996 - 2021 (22) 0.15 Moldova 1996 - 2019 (22) -0.91 Mongolia 2006 - 2020 (15) 0.10 Morocco 1999 - 2021 (23) 1.13 Netherlands 1996 - 2021 (23) 0.51 New Zealand 1996 - 2019 (21) 0.78 Nicaragua 2007 - 2018 (12) 0.08 Niger 1996 - 2019 (24) 0.90 Norway 1981 - 2021 (38) 0.95 Peru 2008 - 2021 (14) 0.76 Philippines 1996 - 2022 (27) 1.76 Poland 1996 - 2021 (25) 0.71 Portugal 1996 - 2022 (24) 0.00 Qatar 2002 - 2017 (14) 1.89 Romania 1996 - 2020 (22) 0.03 Russia 1996 - 2017 (13) -0.19 Saudi Arabia 2003 - 2009 (7) 2.65 Senegal 2015 - 2021 (7) 0.24 Serbia 1998 - 2021 (23) 0.61 Slovakia 1996 - 2022 (24) 0.15 Slovenia 1996 - 2021 (20) 0.28 South Africa 1996 - 2022 (24) 0.28 South Korea 1996 - 2020 (24) 0.63 Spain 1995 - 2021 (25) 0.32 Sweden 1980 - 2021 (37) 0.41 Switzerland 1993 - 2021 (25) 0.69 Tunisia 1996 - 2011 (16) 0.29 Turkey 2010 - 2017 (8) 1.32 USA 1980 - 2021 (38) 0.31 Ukraine 1996 - 2019 (21) 0.17 United Kingdom 1981 - 2021 (36) 0.16 Uruguay 2016 - 2017 (2) 1.16 Uzbekistan 2011 - 2021 (10) 0.21 Vanuatu 2002 - 2007 (6) 0.29 Venezuela 1998 - 2019 (20) 0.55
@@
Note: Thrift theory predicts $\frac{s^{*}}{g(K)}=1$. Free growth theory makes no prediction for these data.
@@
Table 3: Average $\theta_{s}^{*}$ in 92 countries (screen = 0.01). Number of years clearing screen shown in ()